OpenStudy (anonymous):

simplify 3/4i an imaginary number cannot be in the denominator

OpenStudy (anonymous):

what grade is this

OpenStudy (anonymous):

algebra II

OpenStudy (anonymous):

Multiply the top and bottom by "i" That is the factor i/i.

OpenStudy (anonymous):

wouldnt it be a negative i?

OpenStudy (anonymous):

Why do you think you would need negative "i" ?

OpenStudy (anonymous):

because wouldnt you use the discriminant to remove the i?

OpenStudy (anonymous):

There is no discriminant here, at least in the problem as stated. Just try to do that multiplication of top and bottom and see what you get. btw, the support for utilizing the technique I suggested is that any number, complex or not, over the same number is the factor "1" and that is the multiplicative identity so you can legitimately perform that operation.

OpenStudy (anonymous):

okay ill give it a try to that would give me 3i/-1

OpenStudy (anonymous):

oh wait that would get 3i/-4

OpenStudy (anonymous):

\[\frac{ 3 }{ 4i } \times \frac{ i }{ i } = \frac{ 3i }{ 4(i ^{2}) } = \frac{ 3i }{ (4)(-1) } = \frac{ -3i }{ 4 }\]

OpenStudy (anonymous):

Yes, you got it! Good job!

OpenStudy (anonymous):

okay its making it clear now thanks for the help!!

OpenStudy (anonymous):

uw! Good luck to you in all of your studies and thx for the recognition! @TheAnon16

OpenStudy (anonymous):

no thank you for the help ive been stuck on this for a day or two now

OpenStudy (anonymous):

I'm sorry tat was the case. Just make it over to Openstudy more quickly to save yourself some unnecessary frustration.